# Dynamical black hole entropy beyond general relativity from the Einstein frame

### v2

@@ Appendix A The modified canonical energy flux for $f(R)$ gravity
In this appendix we calculate the modified canonical energy flux for $f(R)$ gravity using the covariant phase space formalism. To simplify the calculations, we shall follow the gauge conditions of Hollands:2024vbe that take $\xi^{a}$ to be fixed under the variation. And the gauge conditions on $\delta g_{ab}$ at the event horizon reads:
@@
which are less relaxed than the gauge conditions in Sec.2.
@@ A.1 The covariant phase space formalism and the fundamental identity
where the notation $\cdot$ denotes the contraction of a vector field with the first index of a differential form. It was shown that the Noether current can also be written in the formIyer:1995kg ; Seifert:2006kv
@@
Here, the $(n-2)$-form $\boldsymbol{Q}$ is referred to as the “Noether charge”Iyer:1994ys and the dual vector valued $(n-1)$-form $\boldsymbol{C}_{a}$ vanishes when the equations of motion are satisfied. Taking the first variation of (92) (taking the vector field $\chi^{a}$ to be fixed) and using (90) and (91), we obtain
@@
Considering the variation of (93), we obtain the fundamental identity according toHollands:2012sf
$\boldsymbol{\omega}(\phi;\delta\phi,\mathcal{L}_{\chi}\phi)=\chi\cdot\left[\boldsymbol{E}(\phi)\delta\phi\right]+\chi^{a}\delta\boldsymbol{C}_{a}(\phi)+\mathrm{d}\left[\delta\boldsymbol{Q}[\chi]-\chi\cdot\boldsymbol{\theta}(\phi,\delta\phi)\right].$ (95)
For the case where $\chi^{a}$ is a Killing field of the background $\phi$ such that $\mathcal{L}_{\chi}\phi=0$, we vary the fundamental identity (95) to yield [^6] Note that here the symbol $\mathcal{L}_{\chi}\delta$ is seen as “one” variation, which means that $\boldsymbol{\omega}(\phi;\delta\phi,\mathcal{L}_{\chi}\delta\phi)=\delta\boldsymbol{\theta}(\phi,\mathcal{L}_{\chi}\delta\phi)-(\mathcal{L}_{\chi}\delta)\boldsymbol{\theta}(\phi,\delta_{1}\phi).$ (96) For instance, writing $\boldsymbol{\theta}(\phi,\delta\phi)=\boldsymbol{D}\delta\phi$ formally, where $\boldsymbol{D}$ is a linear operator valued $(n-1)$-form acting on $\delta\phi$ constructed out of $\phi$, its derivatives, and the covariant derivative with respect to the stationary metric $g_{ab}$. If we see $\mathcal{L}_{\chi}\delta$ as “one” variation, we will get $(\mathcal{L}_{\chi}\delta)\boldsymbol{\theta}(\phi,\delta\phi)=(\mathcal{L}_{\chi}\delta\boldsymbol{D})\delta\phi+\boldsymbol{D}(\mathcal{L}_{\chi}\delta^{2}\phi),$ (97) and if we see $\mathcal{L}_{\chi}\delta$ as two subsequent variation, we will get $\mathcal{L}_{\chi}(\delta\boldsymbol{\theta}(\phi,\delta\phi))=(\mathcal{L}_{\chi}\delta\boldsymbol{D})\delta\phi+\delta\boldsymbol{D}(\mathcal{L}_{\chi}\delta\phi)+\boldsymbol{D}(\mathcal{L}_{\chi}\delta^{2}\phi),$ (98) where $\delta\boldsymbol{D}$ constructed out of $\phi$, $\delta\phi$, their derivatives, and the variation of the Christoffel symbols.
$\boldsymbol{\omega}(\phi;\delta\phi,\mathcal{L}_{\chi}\delta\phi)=\chi\cdot\left[\delta\boldsymbol{E}(\phi)\delta\phi\right]+\chi\cdot\left[\boldsymbol{E}(\phi)\delta^{2}\phi\right]+\chi^{a}\delta^{2}\boldsymbol{C}_{a}(\phi)+\mathrm{d}\left[\delta^{2}\boldsymbol{Q}[\chi]-\chi\cdot\delta\boldsymbol{\theta}(\phi,\delta\phi)\right].$ (99)
@@ A.2 The second-order behavior of $S_{\text{dyn}}$ for vacuum perturbations
Now let’s consider the pure gravity theory in which matter fields are absent in the Lagrangian. For vacuum perturbations, we have no external matter sources, $\delta T_{ab}=0$, and (50) states that there is no change of dynamical black hole entropy with respect to affine time at first order, $\Delta\delta S_{\text{dyn}}=0$. Thus we must go to second order in perturbation theory to obtain the leading-order dynamical behavior of the black hold entropy. By (99) and setting $\chi^{a}=\xi^{a}$, in the case that the vacuum equations of motion hold, we obtainHollands:2024vbe
$\boldsymbol{\omega}(g;\delta g,\mathcal{L}_{\xi}\delta g)+\mathrm{d}\left[\xi\cdot\delta\boldsymbol{\theta}(g,\delta g)-\xi\cdot\delta^{2}\boldsymbol{B}_{\mathcal{H}}(g)\right]=\mathrm{d}\left[\delta^{2}\boldsymbol{Q}[\xi]-\xi\cdot\delta^{2}\boldsymbol{B}_{\mathcal{H}}(g)\right]=\frac{\kappa}{2\pi}\mathrm{d}\delta^{2}\boldsymbol{S}_{\text{dyn}}(g),$ (100)
where $\boldsymbol{B}_{\mathcal{H}}$ satisfying $\boldsymbol{\theta}\overset{\mathcal{H}^{+}}{=}\delta\boldsymbol{B}_{\mathcal{H}}$ is defined on $\mathcal{H}$ according to the theorem 1 of Hollands:2024vbe , and it has the following form
@@
where the tensors $\tilde{T}_{(i)}^{b_{1}\cdots b_{i}cd}=\tilde{T}_{(i)}^{(b_{1}\cdots b_{i})(cd)}$ are smooth on $\mathcal{H}$ and are locally and covariantly constructed from the metric, curvature, covariant derivatives of the curvature as well as $\xi^{a}$ and $N^{a}$, with $\xi^{a}$ and $N^{a}$ appearing only algebraically [^7] The vector field $N^{a}$ is introduced byHollands:2024vbe $\nabla_{a}\xi_{b}\overset{\mathcal{H}^{+}}{=}2\kappa N_{[a}\xi_{b]},\quad N^{a}N_{a}\overset{\mathcal{H}^{+}}{=}0,\quad N^{a}\xi_{a}\overset{\mathcal{H}^{+}}{=}1.$ (102) . And the entropy $(n-2)$-form $\boldsymbol{S}_{\text{dyn}}$ is defined on $\mathcal{H}$ by
@@
Next we are going to show that $\boldsymbol{e}_{G}$ is quadratic in $\delta g_{ab}$ and does not depend on $\delta^{2}g_{ab}$. Again, we formally write $\boldsymbol{\theta}(g,\delta g)=\boldsymbol{D}\delta g$ and $\boldsymbol{B}_{\mathcal{H}}(g)=\boldsymbol{C}\mathcal{L}_{\xi}g$ by (101), then $\boldsymbol{\theta}\overset{\mathcal{H}^{+}}{=}\delta\boldsymbol{B}_{\mathcal{H}}$ implies that $\boldsymbol{D}\overset{\mathcal{H}^{+}}{=}\boldsymbol{C}\mathcal{L}_{\xi}$. Thus
@@
Thus we have shown that $\boldsymbol{e}_{G}$ actually does not depend on $\delta^{2}g_{ab}$, and to calculate $\boldsymbol{e}_{G}$ in practice, we need only to calculate $\delta\boldsymbol{\theta}(g;\mathcal{L}_{\xi}\delta g)$ and $\delta^{2}\boldsymbol{B}_{\mathcal{H}}$ and take their parts which are quadratic in $\delta g_{ab}$ by comparing (106) and (108).
@@
$E_{ab}=\frac{1}{16\pi G_{D}}\left(\frac{1}{2}g_{ab}f-\frac{\partial f}{\partial R}R_{ab}+\nabla_{a}\nabla_{b}\frac{\partial f}{\partial R}-g_{ab}\square\frac{\partial f}{\partial R}\right)=0,$ (110)
@@
Suppose that at the late time limit, the spacetime returns to stationary, i.e. $\lim_{v\rightarrow\infty}(f^{\prime}\delta\theta+f^{\prime\prime}\mathcal{L}_{k}\delta R)=0$, then with the above linearized equation of motion, we get
@@
where $\tau$ is the Killing parameter for the null Killing generator $\xi^{a}$, and the Killing parameter $\tau$ is related to the affine parameter $v$ by $v=\frac{C}{\kappa}e^{\kappa\tau}$, which implies that $\mathrm{d}v=\kappa v\mathrm{d}\tau$. (119) gives the same result as (64), and this provides a non-trivial check for our conclusion obtained via the Einstein frame.

### v4

@@ Appendix A The modified canonical energy flux for $f(R)$ gravity
In this appendix we calculate the modified canonical energy flux for $f(R)$ gravity using the covariant phase space formalism. To simplify the calculations, we shall follow the gauge conditions of [1] that take $\xi^{a}$ to be fixed under the variation. And the gauge conditions on $\delta g_{ab}$ at the event horizon reads:
@@
which are less relaxed than the gauge conditions in Sec. II.
@@ A.1 The covariant phase space formalism and the fundamental identity
where the notation $\cdot$ denotes the contraction of a vector field with the first index of a differential form. It was shown that the Noether current can also be written in the form[30, 31]
@@
Here, the $(n-2)$-form $\boldsymbol{Q}$ is referred to as the “Noether charge”[6] and the dual vector valued $(n-1)$-form $\boldsymbol{C}_{a}$ vanishes when the equations of motion are satisfied. Taking the first variation of (93) (taking the vector field $\chi^{a}$ to be fixed) and using (91) and (92), we obtain
@@
Considering the variation of (94), we obtain the fundamental identity according to[18]
$\boldsymbol{\omega}(\phi;\delta\phi,\mathcal{L}_{\chi}\phi)=\chi\cdot\left[\boldsymbol{E}(\phi)\delta\phi\right]+\chi^{a}\delta\boldsymbol{C}_{a}(\phi)\mathrm{d}\left[\delta\boldsymbol{Q}[\chi]-\chi\cdot\boldsymbol{\theta}(\phi,\delta\phi)\right].$ (96)
For the case where $\chi^{a}$ is a Killing field of the background $\phi$ such that $\mathcal{L}_{\chi}\phi=0$, we vary the fundamental identity (96) to yield [^5] Note that here the symbol $\mathcal{L}_{\chi}\delta$ is seen as “one” variation, which means that $\boldsymbol{\omega}(\phi;\delta\phi,\mathcal{L}_{\chi}\delta\phi)=\delta\boldsymbol{\theta}(\phi,\mathcal{L}_{\chi}\delta\phi)-(\mathcal{L}_{\chi}\delta)\boldsymbol{\theta}(\phi,\delta\phi).$ (97) For instance, we write $\boldsymbol{\theta}(\phi,\delta\phi)=\boldsymbol{D}\delta\phi$ formally, where $\boldsymbol{D}$ is a linear operator valued $(n-1)$-form acting on $\delta\phi$ constructed out of $\phi$, its derivatives, and the covariant derivative with respect to the stationary metric $g_{ab}$. If we view $\mathcal{L}_{\chi}\delta$ as “one” variation, we will get $(\mathcal{L}_{\chi}\delta)\boldsymbol{\theta}(\phi,\delta\phi)=(\mathcal{L}_{\chi}\delta\boldsymbol{D})\delta\phi+\boldsymbol{D}(\mathcal{L}_{\chi}\delta^{2}\phi),$ (98) and if we see $\mathcal{L}_{\chi}\delta$ as two subsequent variations, we will get $\mathcal{L}_{\chi}(\delta\boldsymbol{\theta}(\phi,\delta\phi))=(\mathcal{L}_{\chi}\delta\boldsymbol{D})\delta\phi+\delta\boldsymbol{D}(\mathcal{L}_{\chi}\delta\phi)+\boldsymbol{D}(\mathcal{L}_{\chi}\delta^{2}\phi),$ (99) where $\delta\boldsymbol{D}$ is constructed out of $\phi$, $\delta\phi$, their derivatives, and the variation of the Christoffel symbols.
$\begin{split}\boldsymbol{\omega}(\phi;\delta\phi,\mathcal{L}_{\chi}\delta\phi)&=\chi\cdot\left[\delta\boldsymbol{E}(\phi)\delta\phi\right]+\chi\cdot\left[\boldsymbol{E}(\phi)\delta^{2}\phi\right]\\ &+\chi^{a}\delta^{2}\boldsymbol{C}_{a}(\phi)+\mathrm{d}\left[\delta^{2}\boldsymbol{Q}[\chi]-\chi\cdot\delta\boldsymbol{\theta}(\phi,\delta\phi)\right].\end{split}$ (100)
@@ A.2 The second-order behavior of $S_{\text{dyn}}$ for vacuum perturbations
Now let’s consider the pure gravity theory in which matter fields are absent in the Lagrangian. For vacuum perturbations, we have no external matter sources, $\delta T_{ab}=0$, and (50) states that there is no change of dynamical black hole entropy with respect to affine time at first order, $\Delta\delta S_{\text{dyn}}=0$. Thus we must go to second order in perturbation theory to obtain the leading-order dynamical behavior of the black hold entropy. By (100) and setting $\chi^{a}=\xi^{a}$, in the case that the vacuum equations of motion hold, we obtain[1]
$\begin{split}&\boldsymbol{\omega}(g;\delta g,\mathcal{L}_{\xi}\delta g)+\mathrm{d}\left[\xi\cdot\delta\boldsymbol{\theta}(g,\delta g)-\xi\cdot\delta^{2}\boldsymbol{B}_{\mathcal{H}}(g)\right]\\ &=\mathrm{d}\left[\delta^{2}\boldsymbol{Q}[\xi]-\xi\cdot\delta^{2}\boldsymbol{B}_{\mathcal{H}}(g)\right]=\frac{\kappa}{2\pi}\mathrm{d}\delta^{2}\boldsymbol{S}_{\text{dyn}}(g),\end{split}$ (101)
where $\boldsymbol{B}_{\mathcal{H}}$ satisfying $\boldsymbol{\theta}\overset{\mathcal{H}^{+}}{=}\delta\boldsymbol{B}_{\mathcal{H}}$ is defined on $\mathcal{H}$ according to the theorem 1 of [1], and it has the following form
@@
where the tensors $\tilde{T}_{(i)}^{b_{1}\cdots b_{i}cd}=\tilde{T}_{(i)}^{(b_{1}\cdots b_{i})(cd)}$ are smooth on $\mathcal{H}$ and are locally and covariantly constructed from the metric, curvature, covariant derivatives of the curvature as well as $\xi^{a}$ and $N^{a}$, with $\xi^{a}$ and $N^{a}$ appearing only algebraically [^6] The vector field $N^{a}$ is introduced by[1] $\nabla_{a}\xi_{b}\overset{\mathcal{H}^{+}}{=}2\kappa N_{[a}\xi_{b]},\quad N^{a}N_{a}\overset{\mathcal{H}^{+}}{=}0,\quad N^{a}\xi_{a}\overset{\mathcal{H}^{+}}{=}1.$ (103) . And the entropy $(n-2)$-form $\boldsymbol{S}_{\text{dyn}}$ is defined on $\mathcal{H}$ by
@@
Next we are going to show that $\boldsymbol{e}_{G}$ is quadratic in $\delta g_{ab}$ and does not depend on $\delta^{2}g_{ab}$. Again, we formally write $\boldsymbol{\theta}(g,\delta g)=\boldsymbol{D}\delta g$ and $\boldsymbol{B}_{\mathcal{H}}(g)=\boldsymbol{C}\mathcal{L}_{\xi}g$ according to (102), then $\boldsymbol{\theta}\overset{\mathcal{H}^{+}}{=}\delta\boldsymbol{B}_{\mathcal{H}}$ implies that $\boldsymbol{D}\overset{\mathcal{H}^{+}}{=}\boldsymbol{C}\mathcal{L}_{\xi}$. Thus
@@
Thus $\boldsymbol{e}_{G}$ actually does not depend on $\delta^{2}g_{ab}$, and to calculate $\boldsymbol{e}_{G}$ in practice, we need only to calculate $\delta\boldsymbol{\theta}(g;\mathcal{L}_{\xi}\delta g)$ and $\delta^{2}\boldsymbol{B}_{\mathcal{H}}$ and take their parts which are quadratic in $\delta g_{ab}$.
@@
$E_{ab}=\frac{1}{16\pi G_{D}}\left(\frac{1}{2}g_{ab}f-\frac{\partial f}{\partial R}R_{ab}\right.+\left.\nabla_{a}\nabla_{b}\frac{\partial f}{\partial R}-g_{ab}\square\frac{\partial f}{\partial R}\right)=0,$ (111)
@@
Suppose that at the late time limit, the spacetime turns to be stationary, i.e. $\lim_{v\rightarrow\infty}(f^{\prime}\delta\theta+f^{\prime\prime}\mathcal{L}_{k}\delta R)=0$, then with the above linearized equation of motion, we get
@@
Where $\tau$ is the Killing parameter for the null Killing generator $\xi^{a}$, and the Killing parameter $\tau$ is related to the affine parameter $v$ by $v=\frac{C}{\kappa}e^{\kappa\tau}$, which implies that $\mathrm{d}v=\kappa v\mathrm{d}\tau$. (120) gives the same result as (65), and this provides a nontrivial check for our conclusion obtained via the Einstein frame.